OpenStudy (anonymous):

Solve the following.... 5x + 4y = -2 3x - 2y = 12

OpenStudy (anonymous):

use elimination or substitution

OpenStudy (anonymous):

i got x=2. first of all if using elim. distribute 2 to 3x-2y=12. so youd have 5x+4y=-2 6x-4y=24 11x=22 x=2

OpenStudy (anonymous):

multiply second equation by 2 then add solve for x then plug into one of equations to gert y

OpenStudy (anonymous):

to find y, plug in x=2 to an equation. 10+4y=-2, 4y=-12 y=-3

OpenStudy (anonymous):

2,-3 is correct

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

duz it help, curlycute?